OpenStudy (anonymous):

A stone thrown from ground level returns to the same level 4 s after. With what speed was the stone thrown? Take g=10ms−2

OpenStudy (anonymous):

Time of Ascend = Time of Descend

OpenStudy (anonymous):

so..it Takes.2 s To Reach the Topmost Point

OpenStudy (anonymous):

As u Know Velocity @ TopMost Point is0 V = u + at 0 = u - gt u = gt = 2* 10 =20m/s

OpenStudy (anonymous):

time (t) = 2s final velocity (v) = 0 m/s g = - acceleration due to gravity (a) = -10m/s^2 therefore initial velocity (u) = ? (v - u)/t = a therefore, u = v - at u = 0 - ((-10)*2) u = - (-20) u = 20 m/s